OpenStudy (anonymous):

whats this factorial 16! / 2! 14 !

OpenStudy (moosey):

14*13*12.....*1

OpenStudy (amistre64):

what do you believe it will be? youve seen them done a few times by now right?

OpenStudy (anonymous):

120..

OpenStudy (moosey):

120, yes

OpenStudy (amistre64):

i assume its gonna be:\[\frac{16!}{2!\ 14!}\]

OpenStudy (moosey):

sorry

OpenStudy (amistre64):

as long as you know what a factorial is defined by; the rest is just grade school division

OpenStudy (amistre64):

the " ! " sign is just another notation used to express a number; it just so happend that the number it defines is equal to the product itself and the numbers below it down to 1

OpenStudy (amistre64):

2! = 2.1 3! = 3.2.1 4! = 4.3.2.1 5! = 5.4.3.2.1 and thats the pattern to follow, start with the number they give you, and multiply it to everything that is under it

OpenStudy (anonymous):

what about the elimination part?

OpenStudy (anonymous):

give me an example please ?

OpenStudy (amistre64):

the elimination part is the "grade school" division part; recall that anything over itself = 1 2/2 = 1 15/15 = 1 ♫/♫ = 1

OpenStudy (amistre64):

and anything times 1 equals itself right?

OpenStudy (anonymous):

right

OpenStudy (amistre64):

lets start by rewriteing these terms into something we can work with; what would we write 16! as? if you follow my examples above that is

OpenStudy (anonymous):

16 x 15 x 14 x 13 x ....... so on & so on till you get to 1

OpenStudy (amistre64):

right; and the 14! is the in the denominator right? im assuming it is; what is 14! written out?

OpenStudy (anonymous):

i did this answer already i have another one to ask you about ..

OpenStudy (amistre64):

ok

OpenStudy (anonymous):

20! 2! 18!

OpenStudy (anonymous):

idk how to do the symbols sorry

OpenStudy (amistre64):

'sok i can read it :)

OpenStudy (amistre64):

another way to write 20! is: 20 x 19 x 18! would you agree or disagree?

OpenStudy (anonymous):

agree

OpenStudy (amistre64):

goos; and also notice that: 2! = 2 x 1 = 2 20 x 19 x 18! 20 x 19 ------------ = ------- ; do you agree? 2! x 18! 2

OpenStudy (amistre64):

goos meant good lol

OpenStudy (anonymous):

agree & 20 x 19= 308 then what do i do ?

OpenStudy (anonymous):

i mean 380 my bad

OpenStudy (amistre64):

we can still simplify before geting all mathical :) 20 x 19 ------- = 10 x 19 , which is simply 190 then right? 2

OpenStudy (anonymous):

right

OpenStudy (amistre64):

380/2 is fine; but i like my numbers small

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

so if i multiply that that will give me my answer ?

OpenStudy (anonymous):

or divide idk which one ?

OpenStudy (amistre64):

we ended up with: \[\frac{20x19}{2}=\frac{380}{2}=190\] or \[\frac{20x19}{2}=\frac{10x19}{1}=190\]

OpenStudy (anonymous):

yay igot it

OpenStudy (amistre64):

the bar indicates division

OpenStudy (anonymous):

(n + 2) ! ------ n!

OpenStudy (amistre64):

well, lets look at it with numbers; say: n=4 (4+2)! 6x5x4! ------ = ------ = 6 x 5 4! 4! 6 = (4+2) 5 = (4+1) so in the end it just means (n+2) x (n+1) n +2 n +1 ----- n^2 +2n n +2 ------------ n^2 +3n +2

OpenStudy (anonymous):

what is ^ ?